# Quantitative polynomial cohomology and applications to $\textrm{L}^{p}$-measure equivalence

### v1

@@ Acknowledgements
The first author was funded by the National Science Center Grant Maestro-13 UMO-2021/42/A/ST1/00306 and by Fondation Sciences Mathématiques de Paris. He wants to thank Piotr Nowak for many useful exchanges. Both authors thank Claudio Llosa Isenrich and Romain Tessera for their interest, as well as discussions and intuitions they shared with us. We are grateful to Gabriel Pallier for his help with precise questions and references, and for agreeing to write an appendix for a future version of this article.
@@ 1.1 Measure equivalence
In what follows we fix a ME-coupling $(\Omega,X_{\Lambda},X_{\Gamma},m)$, as in the definition of measure equivalence. The fundamental domain $X_{\Gamma}$ determines an equivariant map $p_{\Gamma}:\Omega\to\Gamma$ such that $p_{\Gamma}(x)=e_{\Gamma}\iff x\in X_{\Gamma}$ and $p_{\Gamma}(\gamma.x)=\gamma\,p_{\Gamma}(x)$ for all $\gamma\in\Gamma$ and almost every $x\in\Omega$. This map satisfies $p_{\Gamma}(y)^{-1}.y\in X_{\Gamma}$ for almost every $y\in\Omega$. Similarly, the choice of the fundamental domain $X_{\Lambda}$ determines a map $p_{\Lambda}:\Omega\to\Lambda$ such that $p_{\Lambda}(x)=e_{\Lambda}\iff x\in X_{\Lambda}$, $p_{\Lambda}(\lambda.x)=\lambda\,p_{\Lambda}(x)$ and $p_{\Lambda}(x)^{-1}.x\in X_{\Lambda}$ for all $\lambda\in\Lambda$ and almost every $x\in\Omega$.
Since the actions of $\Gamma$ and $\Lambda$ commute, the action of $\Gamma$ on $\Omega$ passes to the quotient $X_{\Lambda}$ and similarly the action of $\Lambda$ on $\Omega$ passes to the quotient to an action on $X_{\Gamma}$. These quotient actions are better understood by introducing cocycles.
@@
[…]
@@
Let $\gamma_{1},\gamma_{2}\in\Gamma$. We successively integrate the cocycle relation on $x\in X_{\Lambda}$, use the triangle inequality and Minkowski’s inequality to obtain:
@@
Our results will often require mutually cobounded $\textrm{L}^{p}$-ME. The following lemma states that if a ME-coupling is cobounded, up to changing one of the two groups by a finite product, we can assume that one of the two fundamental domains is contained in the other, and this without losing any integrability. Similar statements can be found in [shalom2004harmonic, 2.1.7] or in the proof of [DKLMT-hyperbolicity, Claim 4.5].
@@
We close this section by discussing the related notion of orbit equivalence (OE). We will not really use the definition of OE, but we recall that two countable groups $\Lambda$ and $\Gamma$ are OE if and only if there exists a ME-coupling $(\Omega,X_{\Lambda},X_{\Gamma},m)$ between them satisfying $X_{\Lambda}=X_{\Gamma}$. In particular, such a coupling is mutually cobounded. Notice that while Lemma 1.6 allows us to put one fundamental domain inside the other in the situation of a mutually cobounded ME-coupling, we cannot always get a coupling satisfying $X_{\Lambda}=X_{\Gamma}$. Hence $L^{p}$-OE (that is, the existence of a ME-coupling $(\Omega,X_{\Lambda},X_{\Gamma},m)$ satisfying $X_{\Lambda}=X_{\Gamma}$ that is $L^{p}$-integrable) is a relation stronger than mutually cobounded $L^{p}$-ME.
@@ 1.2 Finiteness properties and filling functions
In this section we introduce a combinatorial version of filling functions, these can be seen as a possible way of extending the more classical Dehn function to higher dimensions. Before this, we recall the definition of the finiteness property $F_{d}$, which will be a standing assumption for most of our results. Most of these definitions and properties are presented in [bader-sauer, Section 6.3] and [young-filling-nilpotent, Section 2].
@@
The (combinatorial) filling volume of an integral chain $c\in C_{k-1}(X,\mathbb{Z})$ is defined as:
@@
(see [alonso-wang-pride, Corollary 3] for higher Dehn functions, the argument also applies to the combinatorial filling functions $\mathrm{cFV}^{j}$). Hence for a group $\Gamma$ of type $F_{k+1}$, we may define (up to the relation $\sim$) the combinatorial filling function $\mathrm{cFV}_{\Gamma}^{k}$ of $\Gamma$ as $\mathrm{cFV}_{X}^{k}$ for any (k1)(k-1)-connected simplicial complex $X$ with such an action.
We say that a function $f:\mathbb{N}\to\mathbb{R}_{>0}$ has polynomial growth or is polynomially bounded if there exists $d\in\mathbb{R}_{+}$ such that $f\lesssim t^{d}$. If there exists $d\in\mathbb{R}_{>0}$ such that $f\sim t^{d}$ we will say that $f$ is polynomial. Wenger exhibited examples of (nilpotent) groups whose Dehn function $\delta_{\Gamma}$ is polynomially bounded but not polynomial [wenger-non-polyn-dehn]. In any case, our results will barely be sensitive to the distinction between polynomial and polynomially bounded. If $\mathrm{cFV}_{\Gamma}^{k}$ is polynomially bounded, we denote
@@
The filling functions $\mathrm{cFV}_{\Gamma}^{k+1}$ are related to their homotopical variants, the higher Dehn functions $\delta_{\Gamma}^{k}$ [young-filling-nilpotent, Section 2]. Indeed, for $k\geq 3$, we have $\mathrm{cFV}_{\Gamma}^{k+1}\sim\delta_{\Gamma}^{k}$. For $k=1,2$, the relationship is more subtle. For $k=2$, it is known that $\delta^{2}_{\Gamma}\lesssim\mathrm{cFV}_{\Gamma}^{3}$ and that there exist groups for which this inequality is strict [abrams-brady-dani-young]. For $k=1$, it is conjectured that the opposite behaviour holds, that is $\mathrm{cFV}_{\Gamma}^{2}\lesssim\delta_{\Gamma}:=\delta_{\Gamma}^{1}$ [brady-kropholler-soroko, 2.29]. This is nearly known to be true, we record the following result, which is a particular case of [brady-kropholler-soroko, 2.28] applied to the convex function $t\mapsto t^{d}$.
@@
One can expect $\deg\mathrm{cFV}^{k}_{\Gamma}$ to attain any real value. Indeed, there are examples of groups $\Gamma$ having higher Dehn functions $\delta_{\Gamma}^{k}\sim t^{d}$ where $d$ ranges over a dense subset of $(k+1)/k,\infty) [$brady-bridson-forester-shankar, Corollary E].
@@ 2 A quantitative version of polynomial cohomology
The goal of this section is to prove Theorem C, which can be thought of as a quantitative analogue of [bader-sauer, Proposition 6.14]. We restate the result here.
@@ 2.1 Quantitative polynomial cohomology
Polynomial cohomology is defined by restricting our attention to cochains with a polynomial growth condition. Following [bader-sauer], we introduce a quantitative version of polynomial cohomology in order to take into account the degree of the polynomial in question.
We first recall how to define (continuous) group cohomology as in [borel-wallach, Chapter IX]. Even though we will restrict our attention to discrete groups, we will give these definitions in the locally compact setting.
@@ 2.2 Polynomially decaying algebras
We introduce some preliminary tools that will be necessary for the proof of Theorem C. These tools are analogous to those in [bader-sauer, 6.2].
@@
[bader-sauer, Lemma 6.3] Let $V$ be an isometric representation on a Banach space. For every $N>0$, the inclusion
@@
$q_{N}(\sum_{\overline{\gamma}\in\Gamma^{n+1}}a_{\overline{\gamma}}(\gamma_{0},\ldots,\gamma_{n}))=\sum_{\overline{\gamma}\in\Gamma^{n+1}}|a_{\overline{\gamma}}|(1+|\gamma_{0}|+\ldots+|\gamma_{n}|)^{N}.$
@@ 2.3 Use of combinatorial filling functions
We first construct an integral chain contraction $h_{*}:C_{*}(X,\mathbb{C})\to C_{*+1}(X,\mathbb{C})$ similar to the one in the proof of [bader-sauer, Proposition 6.13] (but we use the more classical filling functions $\mathrm{cFV}_{X}^{j}$ instead of their weighted versions). First define $h_{-1}:\mathbb{C}\to C_{0}(X,\mathbb{C})$ to be the map sending $1$ to some base vertex $v_{0}\in X^{(0)}$. For some $v\in X^{(0)}$ define $h_{0}(v)$ to be a simplicial geodesic from $v$ to $v_{0}$, this map naturally satisfies $\partial h_{0}+h_{-1}\partial=\mathrm{Id}$ and is integral.
@@
We combine ideas of the proof of [bader-sauer, Lemma 6.7] with Proposition 2.5. Fix $N>0$. We first deal with $k=1$. For $\gamma\in\Gamma$, the chain $g_{1}(1,\gamma)$ is just a simplicial geodesic from $\gamma.v_{0}$ to $v_{0}$. Call the vertices appearing in the geodesic $v_{i}\in V^{(0)}$, where $0\leq i\leq n$. We have $w(v_{i})\leq l(\gamma)$ and $n\leq Cl(\gamma))$ so:
@@
Now let $k\geq 2$ and suppose that the inequality has been proven for $i\leq k-1$. Let $(\gamma_{1},\ldots,\gamma_{k})\in\Gamma^{k}$, and denote by $b:=g_{k}(1,\gamma_{1},\ldots,\gamma_{k})$ and by $c:=g_{k-1}\circ\partial(1,\gamma_{1},\ldots,\gamma_{k})$. Since $g_{*}$ is a chain map, we have that $\partial b=c$ and by definition of $h_{k-1}$ we also have that $b||_{1}=\mathrm{cFV}^{k}(c)$. We now mimic the estimates in the proof of [bader-sauer, Lemma 6.7] to obtain:
@@ 2.4 Comparing quantitative polynomial and ordinary cohomology
We can now prove our quantitative analogue of [bader-sauer, Proposition 6.14].
@@ 2.5 Examples and consequences
Another example of interest are hyperbolic groups, since all of their homological filling functions are linear [lang-hyp-filling-functions, Theorem 4], we obtain the following result.
@@
Since all the homological filling functions of $\Gamma$ are linear [lang-hyp-filling-functions, Theorem 4], we have $N_{k}=\prod_{j=2}^{k}\deg(\mathrm{cFV_{X}^{j}})=1$. We can compute $\alpha_{j}(1)=\alpha_{j-1}(1)+1$ and obtain $\alpha_{j}(1)=j+1$. inductively. Hence $M_{k-1}=\alpha_{k-1}(N_{k})=\alpha_{k-1}(1)=k$. ∎
@@
In particular, if $\Gamma$ is a $CAT(0)$ group acting properly and cocompactly on some $CAT(0)$-space $X$, for $k\geq 2$ we have $N_{k}=\prod_{j=2}^{k}\deg\mathrm{cFV}^{j}_{X}=\prod_{j=2}^{k}\frac{j}{j-1}=k$ [wenger-short-proof], so we have $\alpha_{k-1}(N_{k})=k^{2}+(k-2)k=2k^{2}-2k$. The same computation applies to any non-cocompact arithmetic lattice $\Gamma$ in a semisimple Lie group $G$ with $\mathrm{rank}_{\mathbb{R}}G\geq 3$, for $k<\mathrm{rank}_{\mathbb{R}}G$ [leuzinger-young].
@@ 3.2 Induction of quantitative polynomial cohomology
The fact that $Ib$ is $\Gamma$-equivariant, and that the resulting map $I$ commutes with differentials are standard [shalom2004harmonic, Section 3.2]. The continuity of $I$ in the topology of uniform convergence on compact subsets (which here is just the topology of pointwise convergence) follows from the dominated convergence theorem.
@@ 3.3 Transfer operator and quantitative polynomial cohomology
In this section we define a transfer operator as in [shalom2004harmonic, 3.2.1] for polynomial cohomology and reduced polynomial cohomology for ME-couplings with sufficient integrability.

### v2

@@ Acknowledgements
The first author was funded by the National Science Center Grant Maestro-13 UMO-2021/42/A/ST1/00306 and by Fondation Sciences Mathématiques de Paris. He wants to thank Piotr Nowak for many useful exchanges. Both authors thank Claudio Llosa Isenrich and Romain Tessera for their interest, as well as discussions and intuitions they shared with us. We thank Francesco Fournier-Facio and Roman Sauer for pointing us to their article [18], which motivated us to include a proof of Corollary 0.3 in this new version. Finally, we are indebted to Gabriel Pallier for his generous help and precise references regarding polynomiality of filling functions of nilpotent groups.
@@ 1.1 Measure equivalence
In what follows we fix a ME-coupling $(\Omega,X_{\Lambda},X_{\Gamma},m)$ between $\Lambda$ and $\Gamma$, as in the definition of measure equivalence. The fundamental domain $X_{\Gamma}$ determines an $\Gamma$-equivariant map $p_{\Gamma}:\Omega\to\Gamma$ such that $p_{\Gamma}(x)=e_{\Gamma}\iff x\in X_{\Gamma}$ and $p_{\Gamma}(\gamma.x)=\gamma\,p_{\Gamma}(x)$ for all $\gamma\in\Gamma$ and almost every $x\in\Omega$. This map satisfies $p_{\Gamma}(y)^{-1}.y\in X_{\Gamma}$ for almost every $y\in\Omega$. Similarly, the choice of the fundamental domain $X_{\Lambda}$ determines a $\Lambda$-equivariant map $p_{\Lambda}:\Omega\to\Lambda$ such that $p_{\Lambda}(x)=e_{\Lambda}\iff x\in X_{\Lambda}$, $p_{\Lambda}(\lambda.x)=\lambda\,p_{\Lambda}(x)$ and $p_{\Lambda}(x)^{-1}.x\in X_{\Lambda}$ for all $\lambda\in\Lambda$ and almost every $x\in\Omega$.
Since the actions of $\Lambda$ and $\Gamma$ on $\Omega$ commute, we can define the induced action of $\Gamma$ on $\Omega/\Lambda$ (resp. the induced action of $\Lambda$ on $\Omega/\Gamma$). We will identify these quotient spaces with the corresponding $\Lambda$ and $\Gamma$-fundamental domains. In order to understand the induced actions under this identification, we define cocycles.
@@
Let $(\Gamma,S_{\Gamma})$ and $(\Lambda,S_{\Lambda})$ be two finitely generated measure equivalent groups, with ME-coupling $(\Omega,X_{\Gamma},X_{\Lambda},m)$ and associated cocycles $c_{\Gamma}:\Lambda\times X_{\Gamma}\to\Gamma$ and $c_{\Lambda}:\Gamma\times X_{\Lambda}\to\Lambda$. Denote by $|\cdot|_{S_{\Gamma}}$ and $|\cdot|_{S_{\Lambda}}$ the word metrics associated to the symmetric generating sets $S_{\Gamma}$ and $S_{\Lambda}$. Let $p>0$. We say that the ME-coupling is an $\textrm{L}^{p}$-ME coupling on $\Lambda$ if for all $\gamma\in\Gamma$ we have:
@@
Let $\gamma_{1},\gamma_{2}\in\Gamma$. We successively integrate the cocycle relation on $x\in X_{\Lambda}$ and use the triangle inequality and Minkowski’s inequality to obtain:
@@
Our results will often require mutually cobounded $\textrm{L}^{p}$-ME. The following lemma states that if a ME-coupling is cobounded, up to changing one of the two groups by a finite product, we can assume that one of the two fundamental domains is contained in the other, and this without losing any integrability. Similar statements can be found in [42, 2.1.7] or in the proof of [16, Claim 4.5].
@@
We close this section by discussing the related notion of orbit equivalence (OE). We will not really use the definition of OE, but we recall that two countable groups $\Lambda$ and $\Gamma$ are OE if and only if there exists a ME-coupling $(\Omega,X_{\Lambda},X_{\Gamma},m)$ between them satisfying $X_{\Lambda}=X_{\Gamma}$. In particular, such a coupling is mutually cobounded. Notice that while Lemma 1.6 allows us to put one fundamental domain inside the other in the situation of a mutually cobounded ME-coupling, we cannot always get a coupling satisfying $X_{\Lambda}=X_{\Gamma}$. Hence $\mathrm{L}^{p}$-OE (that is, the existence of a ME-coupling $(\Omega,X_{\Lambda},X_{\Gamma},m)$ satisfying $X_{\Lambda}=X_{\Gamma}$ that is $\mathrm{L}^{p}$-integrable) is a relation stronger than mutually cobounded $\mathrm{L}^{p}$-ME.
@@ 1.2 Finiteness properties and filling functions
In this section we introduce a combinatorial version of filling functions, these can be seen as a possible way of extending the more classical Dehn function to higher dimensions. Before this, we recall the definition of the finiteness property $F_{d}$, which will be a standing assumption for most of our results. Most of these definitions and properties are presented in [5, Section 6.3] and [47, Section 2].
@@
The (combinatorial) filling volume of an integral cycle $c\in Z_{k-1}(X,\mathbb{Z})$ is defined as:
@@
(see [2, Corollary 3] for higher Dehn functions, the argument also applies to the combinatorial filling functions $\mathrm{cFV}^{j}$). Hence for a group $\Gamma$ of type $F_{k+1}$, we may define (up to the relation $\sim$) the combinatorial filling function $\mathrm{cFV}_{\Gamma}^{k}$ of $\Gamma$ as $\mathrm{cFV}_{X}^{k}$ for any (k1)(k-1)-connected simplicial complex $X$ with such an action.
We say that a function $f:\mathbb{N}\to\mathbb{R}_{>0}$ has polynomial growth or is polynomially bounded if there exists $d\in\mathbb{R}_{+}$ such that $f\lesssim t^{d}$. If there exists $d\in\mathbb{R}_{>0}$ such that $f\sim t^{d}$ we will say that $f$ is polynomial. Wenger exhibited examples of (nilpotent) groups whose Dehn function $\delta_{\Gamma}$ is polynomially bounded but not polynomial [45]. In any case, our results will barely be sensitive to the distinction between polynomial and polynomially bounded. If $\mathrm{cFV}_{\Gamma}^{k}$ is polynomially bounded, we denote
@@
The filling functions $\mathrm{cFV}_{\Gamma}^{k+1}$ are related to their homotopical variants, the higher Dehn functions $\delta_{\Gamma}^{k}$ [47, Section 2]. Indeed, for $k\geq 3$, we have $\mathrm{cFV}_{\Gamma}^{k+1}\sim\delta_{\Gamma}^{k}$. For $k=1,2$, the relationship is more subtle. For $k=2$, it is known that $\delta^{2}_{\Gamma}\lesssim\mathrm{cFV}_{\Gamma}^{3}$ and that there exist groups for which this inequality is strict [1]. For $k=1$, it is conjectured that the opposite behaviour holds, that is $\mathrm{cFV}_{\Gamma}^{2}\lesssim\delta_{\Gamma}:=\delta_{\Gamma}^{1}$ [11, 2.29]. This is nearly known to be true, we record the following result, which is a particular case of [11, 2.28] applied to the convex function $t\mapsto t^{d}$.
@@
One can expect $\deg\mathrm{cFV}^{k}_{\Gamma}$ to attain any real value. Indeed, there are examples of groups $\Gamma$ having higher Dehn functions $\delta_{\Gamma}^{k}\sim t^{d}$ where $d$ ranges over a dense subset of $(k+1)/k,\infty) [$10, Corollary E].
@@ 2 A quantitative version of polynomial cohomology
The goal of this section is to prove Theorem C, which can be thought of as a quantitative analogue of [5, Proposition 6.14]. We restate the result here.
@@ 2.1 Quantitative polynomial cohomology
Polynomial cohomology is defined by restricting our attention to cochains with a polynomial growth condition. Following [5], we introduce a quantitative version of polynomial cohomology in order to take into account the degree of the polynomial in question.
We first recall how to define (continuous) group cohomology as in [7, Chapter IX]. Even though we will restrict our attention to discrete groups, we will give these definitions in the locally compact setting.
@@ 2.2 Polynomially decaying algebras
We introduce some preliminary tools that will be necessary for the proof of Theorem C. These tools are analogous to those in [5, 6.2].
@@
[5, Lemma 6.3] Let $V$ be an isometric representation on a Banach space. For every $N>0$, the inclusion
@@
$q_{N}\left(\sum_{\overline{\gamma}\in\Gamma^{n+1}}a_{\overline{\gamma}}(\gamma_{0},\ldots,\gamma_{n})\right)=\sum_{\overline{\gamma}\in\Gamma^{n+1}}|a_{\overline{\gamma}}|(1+|\gamma_{0}|+\ldots+|\gamma_{n}|)^{N}.$
@@ 2.3 Use of combinatorial filling functions
We first construct an integral chain contraction $h_{*}:C_{*}(X,\mathbb{C})\to C_{*+1}(X,\mathbb{C})$ similar to the one in the proof of [5, Proposition 6.13] (but we use the more classical filling functions $\mathrm{cFV}_{X}^{j}$ instead of their weighted versions). First define $h_{-1}:\mathbb{C}\to C_{0}(X,\mathbb{C})$ to be the map sending $1$ to some base vertex $v_{0}\in X^{(0)}$. For some $v\in X^{(0)}$ define $h_{0}(v)$ to be a simplicial geodesic from $v$ to $v_{0}$, this map naturally satisfies $\partial h_{0}+h_{-1}\partial=\mathrm{Id}$ and is integral.
@@
We combine ideas of the proof of [5, Lemma 6.7] with Proposition 2.5. Fix $N>0$. We first deal with $k=1$. For $\gamma\in\Gamma$, the chain $g_{1}(1,\gamma)$ is just a simplicial geodesic from $\gamma.v_{0}$ to $v_{0}$. Call the vertices appearing in the geodesic $v_{i}\in V^{(0)}$, where $0\leq i\leq n$. We have $w(v_{i})\leq l(\gamma)$ and $n\leq Cl(\gamma))$ so:
@@
Now let $k\geq 2$ and suppose that the inequality has been proven for $i\leq k-1$. Let $(\gamma_{1},\ldots,\gamma_{k})\in\Gamma^{k}$, and denote by $b:=g_{k}(1,\gamma_{1},\ldots,\gamma_{k})$ and by $c:=g_{k-1}\circ\partial(1,\gamma_{1},\ldots,\gamma_{k})$. Since $g_{*}$ is a chain map, we have that $\partial b=c$ and by definition of $h_{k-1}$ we also have that $b||_{1}=\mathrm{cFV}^{k}(c)$. We now mimic the estimates in the proof of [5, Lemma 6.7] to obtain:
@@ 2.4 Comparing quantitative polynomial and ordinary cohomology
We can now prove our quantitative analogue of [5, Proposition 6.14].
@@ 2.5 Examples and consequences
Another example of interest are hyperbolic groups, since all of their homological filling functions are linear [30, Theorem 4], we obtain the following result.
@@
Since all the homological filling functions of $\Gamma$ are linear [30, Theorem 4], we have $N_{k}=\prod_{j=2}^{k}\deg(\mathrm{cFV_{X}^{j}})=1$. We can compute $\alpha_{j}(1)=\alpha_{j-1}(1)+1$ and obtain $\alpha_{j}(1)=j+1$. inductively. Hence $M_{k-1}=\alpha_{k-1}(N_{k})=\alpha_{k-1}(1)=k$. ∎
@@
In particular, if $\Gamma$ is a $CAT(0)$ group acting properly and cocompactly on some $CAT(0)$-space $X$, for $k\geq 2$ we have $N_{k}=\prod_{j=2}^{k}\deg\mathrm{cFV}^{j}_{X}=\prod_{j=2}^{k}\frac{j}{j-1}=k$ [44], so we have $\alpha_{k-1}(N_{k})=k^{2}+(k-2)k=2k^{2}-2k$. The same computation applies to any non-cocompact arithmetic lattice $\Gamma$ in a semisimple Lie group $G$ with $\mathrm{rank}_{\mathbb{R}}G\geq 3$, for $k<\mathrm{rank}_{\mathbb{R}}G$ [31].
@@ 3.2 Induction of quantitative polynomial cohomology
The fact that $Ib$ is $\Gamma$-equivariant, and that the resulting map $I$ commutes with differentials are standard [42, Section 3.2]. The continuity of $I$ in the topology of uniform convergence on compact subsets (which here is just the topology of pointwise convergence) follows from the dominated convergence theorem.
@@ 3.3 Transfer operator and quantitative polynomial cohomology
In this section we define a transfer operator as in [42, 3.2.1] for polynomial cohomology and reduced polynomial cohomology for ME-couplings with sufficient integrability.